# Factor Graph Optimization of Error-Correcting Codes for Belief Propagation Decoding

### v1

@@ 4       H = 1-2*H
7             Loss(H,x,y).backward()
8             lambdas = H/H.grad+eps
9             lambdas = sorted(lambdas[lambdas>0].view(-1))[:110]
10             H = H - H.grad*lambdas[argmin([Loss(bin(H-lambda*H.grad),x,y) for lambda in lambdas]]
@@ 5 Experiments
We evaluate our framework on five classes of linear codes: various Low-Density Parity Check (LDPC) codes [23, 1], Polar codes [2], Reed Solomon codes [55], Bose–Chaudhuri–Hocquenghem (BCH) codes [8] and random codes. All the parity check matrices are taken from [25] except the LDPC codes created using the popular Progressive Edge Growth framework [28, 41].
We consider three types of channel noise under BPSK modulation. We first test our framework with the canonical AWGN channel given as $y=c_{s}+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,\sigma I_{n})$. We also consider the Rayleigh fading channel, where $y=h\odot c_{s}+\varepsilon$, with $h$ the iid Rayleigh distributed fading vector with coefficient $1$ and $\varepsilon$ the regular AWGN noise, where we assume ideal channel state information. Finally, we consider the AWGN channel with bursty noise simulating wireless channel interference as $y=c_{s}+\varepsilon+\zeta$ with $\varepsilon$ the AWGN and $\zeta_{i}\sim\mathcal{N}(0,\sqrt{2}\sigma)$ with probability $\rho=0.1$ and $\zeta_{i}=0$ with probability $1-\rho$.
The results are reported as negative natural logarithm bit error rates (BER) for three different normalized SNR values ($E_{b}/N_{0}$), following the conventional testing benchmark, e.g., [46, 12]. BP-based results are obtained after $L=5$ BP iterations in the first row and $L=15$ in the second row of the results tables. During testing, at least $10^{5}$ random codewords are decoded, to obtain at least $50$ frames with errors at each SNR value. For this section, we performed a small hyperparameter search as reported in Appendix A, where the final code is selected to have the lowest average BER on the SNR test range.
@@
LDPC PGE2(64,32) 4.38 4.38 5.12 5.13 6.04 6.04 4.45 4.44 5.19 5.19 6.10 6.10 4.08 4.08 4.44 4.44 4.81 4.81 4.10 4.10 4.46 4.47 4.85 4.85 4.07 4.06 4.69 4.69 5.43 5.43 4.07 4.06 4.70 4.69 5.43 5.44
LDPC PGE5(64,32) 6.02 6.63 8.20 9.06 10.95 12.30 6.53 7.13 8.73 9.48 11.56 12.20 5.63 6.19 6.86 7.52 8.31 9.02 6.22 6.96 7.48 8.34 8.82 9.85 5.18 5.68 6.97 7.75 9.34 10.19 5.59 6.12 7.41 8.06 9.51 10.13
LDPC PGE10(64,32) 3.98 4.27 5.17 5.77 6.70 7.67 5.56 6.25 7.22 8.28 9.13 10.59 3.52 3.71 4.18 4.47 4.95 5.30 5.02 5.60 6.00 6.72 7.11 7.90 3.48 3.67 4.47 4.90 5.75 6.46 4.26 4.73 5.50 6.22 7.01 8.01
The results are provided in Table 1, where our method means BP applied on the learned code initialized by the given classical code. We also provide in Appendix B the same table with a broader SNR range ($E_{b}/N_{0}\in\{3,\dots,7\})$. We provide the improvement statistics (i.e., mean, std, min, max) in dB on all the sparse codes in Figure 2 and we extend the analysis to all the codes in Appendix F.
@@ 6 Analysis
We provide in Figure 3 the performance of the proposed method on random codes initialized with different sparsity rates. The parity check matrix is initialized in a systematic form $H=[I_{n-k},P]$ for full rank initialization, where $P\sim\text{Bern}^{k\times(n-k)}(p)$. We can observe that the framework can greatly improve the performance of the original random code. Most importantly, we can observe that different initializations provide convergence to different local optima and that better initialization generally induces convergence to a better minimum. Performance on other code lengths is provided in Appendix G.
@@
In Figure 4 we provide the performance of the method on constrained systematic random codes, i.e., $H=[P^{T},I_{n-k}]$ with $P\sim\text{Bern}_{p}^{k\times(n-k)}$. The random codes from paragraph 6 are constrained to maintain their systematic form during the optimization, i.e., only the parity matrix elements of $P$ are optimized. The optimization is performed by backpropagating over the $P$ tensor only, similarly to having a hard structure constraint on the identity part of $H$. While maintaining a structure of interest, we can observe this regularization can further improve the convergence quality (e.g., $p=0.1$) compared to the unconstrained setting of Figure 3. Performance on other code lengths is provided in Appendix G.
@@
In Figure 6 we present the reduction of the sparsity of the codes created by the framework. Here, $\Delta=100({S_{b}-S_{o}})/{S_{b}}$ represents the sparsity ratio, with $S_{b}$ and $S_{o}$ being the sparsity of the baseline code and our code, respectively. We can observe that optimization always provides sparser code. Nevertheless, we also observed that the optimization does not modify the girth of the code.
@@ 7 Conclusions
We present a novel data-driven optimization method of binary linear block codes for the Belief Propagation algorithm. The proposed framework enables the differentiable optimization of the factor graph via weighted tensor representation. The optimization is efficiently carried out via a tailor-made grid search procedure that is aware of the binary constraint of the optimization problem.

### v2

@@ 4       $\Omega$ = 1-2*H
7             Loss(bin($\Omega$),x,y).backward()
8             lambdas = $\Omega$/$\Omega$.grad
9             lambdas = sorted(lambdas[lambdas>0].view(-1))[:50]
10             $\Omega$ = $\Omega$ $\Omega$.grad*lambdas[argmin([Loss(bin($\Omega$lambda*$\Omega$.grad),x,y) for lambda in lambdas]]
@@
Training
The optimization parameters are the following: the initial $H$ (i.e., initial $\Omega$), the maximum number of optimization steps (if convergence is not reached), the number and quality of the data samples, the grid search length, and the number of BP iterations.
We assume that an initial $H$ is given by the user as the code to be improved. The number of optimization steps is set to 20 iterations. The training noise is sampled randomly per batch in the $\{3,\dots,7\}$ normalized SNR (i.e. $E_{b}/N_{0}$) range but can be modified according to the noise setting of interest. The number of data samples per optimization iteration is set to 4.9M for every code as sufficient gradient estimation, and the data samples are required to have non-zero syndrome. Because of computational constraints, the number of BP iterations during training is fixed and set to 5, while other ranges or values of interest can be used instead. For faster optimization, the grid search is heuristically restricted to the first 50 smallest step sizes as the optimal step size is generally in the vicinity of the working point (Appendix C) . Training and experiments are performed on $8\times 12\mathrm{GB}$ GeForce RTX 2080 Ti GPUs and require 2.96 minutes on average per optimization step.
@@ 5 Experiments
LDPC PEG2(64,32) 4.38 4.38 5.12 5.13 6.04 6.04 4.45 4.44 5.19 5.19 6.10 6.10 4.08 4.08 4.44 4.44 4.81 4.81 4.10 4.10 4.46 4.47 4.85 4.85 4.07 4.06 4.69 4.69 5.43 5.43 4.07 4.06 4.70 4.69 5.43 5.44
LDPC PEG5(64,32) 6.02 6.63 8.20 9.06 10.95 12.30 6.53 7.13 8.73 9.48 11.56 12.20 5.63 6.19 6.86 7.52 8.31 9.02 6.22 6.96 7.48 8.34 8.82 9.85 5.18 5.68 6.97 7.75 9.34 10.19 5.59 6.12 7.41 8.06 9.51 10.13
LDPC PEG10(64,32) 3.98 4.27 5.17 5.77 6.70 7.67 5.56 6.25 7.22 8.28 9.13 10.59 3.52 3.71 4.18 4.47 4.95 5.30 5.02 5.60 6.00 6.72 7.11 7.90 3.48 3.67 4.47 4.90 5.75 6.46 4.26 4.73 5.50 6.22 7.01 8.01
We evaluate our framework on five classes of linear codes: various Low-Density Parity Check (LDPC) codes (Gallager, 1962; Abu-Surra et al., 2010), Polar codes (Arikan, 2008), Reed Solomon codes (Reed & Solomon, 1960), Bose–Chaudhuri–Hocquenghem (BCH) codes (Bose & Ray-Chaudhuri, 1960) and random codes. All the parity check matrices are taken from (Helmling et al., 2019) except the LDPC codes created using the popular Progressive Edge Growth framework (Hu et al., 2005; MacKay, ).
We consider three types of channel noise under BPSK modulation. We first test our framework with the canonical AWGN channel given as $y=c_{s}+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,\sigma I_{n})$. We also consider the Rayleigh fading channel, where $y=h\odot c_{s}+\varepsilon$, with $h$ the iid Rayleigh distributed fading vector with coefficient $1$ and $\varepsilon$ the regular AWGN noise, where we assume ideal channel state information. Finally, we consider the AWGN channel with Gaussian mixture channel (also referred to as bursty noise) simulating wireless channel interference as $y=c_{s}+\varepsilon+\zeta$ with $\varepsilon$ the AWGN and $\zeta_{i}\sim\mathcal{N}(0,\sqrt{2}\sigma)$ with probability $\rho=0.1$ and $\zeta_{i}=0$ with probability $1-\rho$.
The results are reported as negative natural logarithm bit error rates (BER) for three different normalized SNR values ($E_{b}/N_{0}$), following the conventional testing benchmark, e.g., (Nachmani & Wolf, 2019; Choukroun & Wolf, 2022a). BP-based results are obtained after $\ell=5$ BP iterations in the first row and $\ell=15$ in the second row of the results tables. During testing, at least $10^{5}$ random codewords are decoded, to obtain at least $50$ frames with errors at each SNR value. For this section, we performed a small hyperparameter search as reported in Appendix A, where the final code is selected to have the lowest average BER on the SNR test range.
The results are provided in Table 1, where our method means BP applied on the learned code initialized by the given classical code. We also provide in Appendix B the same table with a broader SNR range ($E_{b}/N_{0}\in\{3,\dots,7\})$. We provide the improvement statistics (i.e., mean, std, min, max) in dB on all the sparse codes in Figure 2 and we extend the analysis to all the codes in Appendix F. For completeness, we provide in Appendix H a comparison with the genetic algorithm of Elkelesh et al. (2019) where our method demonstrates much better performance while being faster by orders of magnitude.
@@ 6 Analysis
We provide in Figure 3 the performance of the proposed method on random codes initialized with different sparsity rates. The parity check matrix is initialized in a systematic form $H=[I_{n-k},P]$ for full rank initialization, where $P\sim\text{Bern}^{(n-k)\times k}(p)$. We can observe that the framework can greatly improve the performance of the original random code. Most importantly, we can observe that different initializations provide convergence to different local optima and that better initialization generally induces convergence to a better minimum. Performance on other code lengths is provided in Appendix G. Finally, good initialization (i.e., (large) well-performing sparse codes under BP decoding) requires perturbation at initialization or during training (c.f., Appendix A) in order to get extracted from local minima.
@@
In Figure 4 we provide the performance of the method on constrained systematic random codes as described in the previous paragraph, while here, we constrain them to maintain their systematic form during the optimization, i.e., only the parity matrix elements of $P$ are optimized (via $\Omega$). The optimization is performed by backpropagating over the $P$ tensor only, similarly to having a hard structure constraint on the identity part of $H$. While maintaining a structure of interest, we can observe this regularization can further improve the convergence quality (e.g., $p=0.1$) compared to the unconstrained setting of Figure 3. Performance on other code lengths is provided in Appendix G.
@@
In Figure 6 we present the sparsification of the codes created by the framework. Here, $\Delta=100({S_{b}-S_{o}})/{S_{b}}$ represents the sparsity ratio, with $S_{b}$ and $S_{o}$ being the sparsity of the baseline code and our code, respectively. We can observe that optimization always provides sparser code. Nevertheless, we also observed that the optimization does not modify the girth of the code.
@@ 7 Conclusions
We present a novel gradient-based optimization method of binary linear block codes for the Belief Propagation algorithm. The proposed framework enables the differentiable optimization of the factor graph via weighted tensor representation. The optimization is efficiently carried out via a tailor-made grid search procedure that is aware of the binary constraint of the optimization problem.
@@ Appendix G More Random Codes
Appendix H Comparison with Genetic Algorithm
We provide in table 4 a comparison with the genetic algorithm of Elkelesh et al. (2019). We note that the method requires 230 offspring/code evaluations per iteration, with 300 iterations (Fig. 7 in (Elkelesh et al., 2019)) or even an infinite loop (cf. the provided MATLAB code). Our algorithm is tested on 50 line-search steps as described in in the paper on 2 to 25 iterations (cf. App. D), which means that Elkelesh et al. (2019) requires approximately 25 to 313 times more computations than our proposed method. The performance presented are for 75 and 150 iterations of the algorithm, representing around 70 and 140 times slower performance than our approach, respectively, while they remain below our performance. We note here, as described in the paper, that combining the methods by allowing the perturbation of the parity-check matrix at a local minima may allow the discovery of other better local optimum.
Table 4: A comparison of the negative natural logarithm of Bit Error Rate (BER) for several normalized SNR values of our method with classical codes. Higher is better. BP results are provided for 5 iterations in the first row and 15 in the second row. The best results are in bold. The training range is defined as $E_{b}/N_{0}=\{5\}$. GA denotes the genetic algorithm of Elkelesh et al. (2019) with $k$ iterations.
Channel AWGN
Method BP Our GA $k=75$ GA $k=150$
$E_{b}/N_{0}$ 4 5 6 4 5 6 4 5 6 4 5 6
CCSDS(128,64) 6.46 7.32 9.61 10.83 13.99 15.43 7.34 8.61 10.48 12.26 14.37 16.00 6.86 7.87 9.95 11.31 13.38 15.56 7.09 8.23 10.40 11.79 14.08 16.04